SUB-MILLIMETER/TERAHERTZ SPECTROSCOPY OF FeH AND FeD (X^{4}Δ_{i})

\underline{M. P. BUCCHINO} and L. M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, Arizona 85721.

Direct measurements of the lowest pure rotational transitions of the Ω = 7/2, 5/2, and 3/2 ladders of FeD and of the Ω = 1/2 ladder of FeH in the ^{4}Δ_{i} ground electronic states have been conducted using millimeter/sub-millimeter direct absorption techniques in the frequency range 530-810 GHz. Both species were created in an AC discharge from the reaction of Fe(CO)_{5} with H_{2} or D_{2}. In several of these transitions lambda-doubling and proton/deuterium hyperfine structure have been resolved. These data improve on the zero field frequencies predicted from previous LMR data by several MHz.^{a} Accurate frequencies are crucial for sub-millimeter searches for FeH toward the interstellar medium. Searches for additional transitions are in progress.

a J.M. Brown, H. Korsgen, S.P. Beaton, & K.M. Evenson, J. Chem. Phys. 124, 234309 (2006)